# Quantum transistors for heat flux in and out of working substance parts: harmonic vs transmon and Kerr environs

### v1

@@ VI Conclusion
Quantum thermal transistors with Markovian baths have been rigorously studied and worked upon. Non-Markovian systems, however, are often more plausible in realistic situations. In this work, we considered a type of transistor which interacts with the environment via a collisional model, thereby deviating from Markovianity. We found that such an apparatus, which we referred to as a working-substance thermal transistor, does manifest the transistor effect. Such a nomenclature owes its origin to the fact that we focused on the heat currents which flow in and out of each qubit of the working substance due to the heat flow from and to the different parts of the system and environment, instead of on the currents which flow in and out of a single bath. We depicted how the figure of merit of the transistor, i.e. the amplification of the heat currents, vary with the temperature of the modulating environment, system-environment coupling and the interaction time. We have also analyzed the variation of amplification when one of the baths, that is not the modulating one, is removed.
@@
Furthermore, we studied the effects of frail non-linearities in the environment, which are either of transmon- or Kerr-type. We found a significant amplification in each of these cases, and showed that both the type of non-linearities prove to be beneficial in different temperature regimes. In all of these cases, we also analyzed the variation of amplification with time. We found a non-monotonicity in the behavior of amplification versus time, which potentially is a manifestation of the inherent non-Markovianity in the system. We also quantified the amount of non-Markovianity by utilizing the distinguishability-based measure of Breuer, Laine and Piilo. Specifically, we found that the quantity of interest, i.e. the amplification, retains a memory of the non-Markovian interactions for a longer time than what the BLP measure itself is able to detect.
@@ Acknowledgements.
We acknowledge computations performed using Armadillo Sanderson and Curtin (2016, 2018) on the cluster computing facility of Harish-Chandra Research Institute (HRI), India. We acknowledge the use of QIClib – a modern C++ library for general purpose quantum information processing and quantum computing. DB thanks HRI for support during visits.

### v2

@@ VII Conclusion
Quantum thermal transistors with Markovian baths have been rigorously studied and worked upon. Non-Markovian systems, however, are often more plausible in realistic situations. In this work, we considered a type of transistor that interacts with the environment via a collisional model consisting of a rail of qutrits, thereby deviating from Markovianity. The working substance consists of three qubits, i.e. left, right, and middle qubits. We found that such an apparatus, which we referred to as a working-substance thermal transistor, does manifest the transistor effect. The interactions between the right-middle and the left-middle qubits are considered to be equal, but there is no interaction between the left and right qubits. Such a nomenclature owes its origin to the fact that we focused on the heat currents which flow in and out of each qubit of the working substance due to the heat flow from and to the different parts of the system and environment, instead of on the currents which flow in and out of a single bath. We depicted how the figure of merit of the transistor, i.e. the amplification of the heat currents, vary with the temperature of the modulating environment, system-environment coupling and the interaction time. We have also analyzed the variation of amplification when one of the baths, that is not the modulating one, is removed.
@@
Furthermore, we studied the effects of frail non-linearities in the environment, which are either of transmon- or Kerr-type. We found a significant amplification in each of these cases, and showed that both the type of non-linearities prove to be beneficial in different temperature regimes. In all of these cases, we also analyzed the variation of amplification with time. We found a non-monotonicity in the behavior of amplification versus time, which potentially is a manifestation of the inherent non-Markovianity in the system. We also quantified the amount of non-Markovianity by utilizing the distinguishability-based measure of Breuer, Laine and Piilo. Specifically, we found that the quantity of interest, i.e. the amplification, retains a memory of the non-Markovian interactions for a longer time than what the BLP measure itself is able to detect. We further analyze two scenarios. In the first, the couplings among the left-middle, right-middle, and left-right qubit pairs are taken to be unequal, corresponding to the asymmetric case. In the second, all three couplings are chosen to be equal, representing the symmetric case. In both settings, we observe sufficient amplification for both the left and right qubits of the working substance. We also demonstrate a pronounced transistor effect for the left and right qubits of the working substance when the environment consists of qubits, and separately when the environment pertains non-linearities, following the same collisional model. We have also examined the non-Markovianity measure in this scenario.
@@ Acknowledgements.
We acknowledge computations performed using Armadillo Sanderson and Curtin (2016, 2018) on the cluster computing facility of Harish-Chandra Research Institute (HRI), India. We acknowledge the use of QIClib – a modern C++ library for general purpose quantum information processing and quantum computing. DB thanks HRI for support during visits. AB acknowledges support from ‘INFOSYS scholarship for senior students’ at Harish Chandra Research Institute, India.
Appendix
Appendix A Amplification with two qubits and two environments
where $J_{L}$ and $J_{R}$ are the heat currents flowing into the left and right qubits respectively. In this case, we observe that when the energy difference between the ground state and first excited state is the same for left and right qubit, then there is no amplification. However, when we consider unequal energy differences, there is a finite amplification in a small range of $T_{L}$. The range of $T_{L}$ is from 0 to 2.488 when $\omega_{L}=1$ and $\omega_{R}=2$, $T_{R}=4,g=4,\Delta=5$ and $t=1$. In conventional classical transistors, the three components, i.e. emitter, base and collector, are all essential to observe the transistor effect, and removing any one of them suppresses this behavior. Prototypically, a two-terminal electronic device that demonstrates transistor-like effects cannot be regarded as a transistor in the classical sense. However, in the quantum mechanical domain, a two-terminal device that exhibits properties similar to a transistor can be referred to as a quantum transistor. In our work, we observe such a transistor effect in a specific regime of the left-bath temperature, $T_{L}$, when the energy gaps of the first and second excited states differ between the left and right qubits. This effect is purely quantum in nature, being absent in the corresponding classical counterparts.